OpenStudy (anonymous):

HELP, Need an easy way to show a 7th grader how to do this kind of problems!! Ratio of cats to dogs was 3 to 2. If there were 170 cats and dogs in all, how many were dogs.9:37 pm

OpenStudy (saifoo.khan):

7 : 3 170 : x cross multiply 7(x) = 170(3) x = 510 / 7 x = 72.85 x = 73 dogs

OpenStudy (anonymous):

3 170 -= ____ 2 x Then you do 2x170/3=x

OpenStudy (saifoo.khan):

oh sorry sorry sorry!

OpenStudy (anonymous):

Sooo its 3/2 is equal to 170/113

OpenStudy (saifoo.khan):

3 : 2 170 : x cross multiply 3(x) = 170(2) x = 340 / 3 x = 113.3 x = 113 dogs

OpenStudy (anonymous):

thats not a given answer!!!

OpenStudy (anonymous):

is answer 102 cats and 68 dogs?

OpenStudy (anonymous):

68 is one of the answers. how did you get it?

OpenStudy (anonymous):

easy... ratio of cats to dog is 3:2 from total 170 so 3/5 *170 = cats(102) and 2/5 *170 is dogs (68) .. adds up to 170

OpenStudy (anonymous):

ok thank u!